Suppose that x has a binomial distribution with n = 15 and p = 0.3. Is the normal approximation reasonable for this binomial distribution?
Answer by robertb(5830)

It probably won't be.
np = 15*0.3 = 4.5 < 5, and
n(1-p) = 15*0.7 = 10.5 > 5.
Both had to be at least 5.